A recent field experiment examined how the genotype at one locus affected the survival of sunflowers....

  1. A recent field experiment examined how the genotype at one locus affected the survival of sunflowers. 200 individuals with the AA, Aa, and aa genotypes were planted. 8pts

Survival was surveyed at the time of flowering.      

AA           Aa          aa           

Number surviving                   170           20           20          

  1. What are the relative fitnesses for each genotype?
  2. Explain the difference between absolute and relative fitness, and why relative fitness is more useful for explaining evolution.
Answer #1

1. Relative fitness

Survival of each genotype / highest genotype

AA = 170/170 = 1

Aa = 20/170 = 0.12

aa = 20/170 = 0.12

2 .

Absolute fitness is ratio between the number of individuals with that genotype after selection to those before selection. Relative fitness is quantified as the average number of surviving progeny of a particular genotype compared with average number of surviving progeny of competing genotypes after a single generation.

Since it compares in the same generation, relative fitness is very useful to find out and in less time